# Few-shot Species Range Estimation

### v1

@@ 6 Conclusion
We have limited knowledge regarding the geographic distributions of the majority of species on earth. This lack of understanding is further hampered by the fact that we also have insufficient data to train models to estimate their ranges. To address this problem, we introduced FS-SINR, a new approach for few-shot species range estimation. We demonstrated that our approach is able to fuse data from different modalities at inference time to make plausible range predictions for previously unobserved species. Our quantitative analysis, using expert-derived range maps, show a 5-10% performance improvement compared to current approaches in the few-shot setting, i.e., when the number of observations equals ten, for previously unseen species. Additionally, we also outperform existing methods in the zero-shot setting. While our results are promising, they also indicate that there are many open challenges in this important task.
Acknowledgements. We thank the iNaturalist community for making the species observation data available.
@@
Another issue associated with training models on species observation data is that there is a risk that sensitive information (e.g., the locations of protected species) could be leaked or extracted from the models. As a result of this concern, the models in this work were only trained on publicly available information which does not include any sensitive observations. Finally, our approach integrates predictions from pretrained large language models. These models are known to be biased and capable of hallucinating fabricated outputs. Spatially localizing the outputs of such models runs the risk of amplifying such biases if used inappropriately.
Appendix A Implementation Details
A.1 FS-SINR
Below we provide implementation details. Upon acceptance, we will release the code so others can reproduce our results.
A.1.1 Model Architecture
@@
The architecture for FS-SINR consists of four components: the location encoder, $f$; the text encoder, $f$; the image encoder, $a$; the transformer encoder, $e$; and the species decoder, $s$. These components comprise of 8,154,368 learnable parameters in total. All non-linearities in FS-SINR are ReLUs.
@@
The location encoder, $f$, is identical to the the one used in Hamilton et al. (2024), which is taken from  Cole et al. (2023). It is composed of an initial linear layer and ReLU nonlinearity followed by four residual layers, where each is a two layer fully connected network with residual connections (He et al., 2015) between the input and output of each residual layer. Each layer contains 256 neurons, and there are 527,616 learnable parameters in total.
@@
The text encoder, $t$, follows the structure of text-based species encoder from Hamilton et al. (2024). In $t$, a pretrained and frozen large language model, GritLM (Muennighoff et al., 2025), is used to produce a fixed 4,096 length embedding from input text. This is then passed through a smaller network to reduce the dimensionality to 256. This smaller network comprises of two residual layers with a hidden layer size of 512. In total, the text encoder contains 3,410,432 learnable parameters.
@@
The image encoder, $a$, has a similar structure to $t$. In $a$, a pretrained and frozen vision transformer, EVA-02 (Fang et al., 2024), pretrained on images from 10,000 species from the iNaturalist species classification dataset (Van Horn et al., 2021), is used to produce a fixed 1,024 length embedding from an input image, by extracting the CLS token from the final layer of the model. This is then passed through a smaller network to reduce the dimensionality to 256. This smaller network comprises of two residual layers with a hidden layer size of 512. In total, the image encoder contains 1,837,568 learnable parameters. In Table A1 we also show zero-shot results using a DINOv2-large image encoder (Oquab et al., 2024) where all other architecture choices remain the same.
@@
The transformer encoder, $e$, takes in an arbitrary length set of unordered 256 dimensional tokens produced by $f$, $t$, and $a$, as well as two learned tokens that are added to each set of inputs. The CLS, class, token produces the species range, and a “Register” token, inspired by Darcet et al. (2024), acts as an additional repository of global information during encoding. Element-wise addition between each token and one of five learned 256 dimensional “token type embeddings” is performed to allow the model to differentiate between tokens from different sources. The transformer itself is composed of four transformer encoder layers, implemented using PyTorch’s nn.TransformerEncoderLayer (Paszke et al., 2019), based on Vaswani et al. (2017). Key-query-value multi-head attention is used with two “heads”. The feed forward components contain 512 neurons per layer, while the token dimensionality is 256. Layer norm is used in each layer, using a default epsilon value of 1e-5 for enhanced numerical stability. In total, $e$ contains 2,176,256 learnable parameters.
Finally the species decoder, $s$, is a simple fully connected network with two hidden layers. Each layer contains 256 neurons, and in total the decoder contains 197,376 learnable parameters.
A.1.2 Training

### v2

@@ 6 Conclusion
The scientific community has limited knowledge on the geographical distributions of the majority of species on Earth. This lack of understanding is further hampered by the fact that we also have insufficient data to train models to estimate their ranges. To address this problem, we introduced FS-SINR, a new approach for few-shot species range estimation. We demonstrated that FS-SINR is able to fuse data from different modalities at inference time in a feedforward manner to efficiently make plausible range predictions for previously unobserved species. Our quantitative analysis, using expert-derived range maps, shows a 5-10% performance improvement compared to current approaches in the few-shot setting, i.e., when the number of observations equals ten, for previously unseen species. In addition, we also outperform existing methods in the zero-shot setting. While our results are promising, they also indicate that there are many open challenges in this important task.
Acknowledgements. We thank the iNaturalist community for making the species observation data available. OMA was supported by a Royal Society Research Grant. MH and SM were supported by NSF grants 2329927 and 2406687.
@@
Another issue associated with training models on species observation data is that there is a risk that sensitive information (e.g., the locations of protected species) could be leaked or extracted from the models. To respect this concern, the models in this work were trained using only publicly available information which does not include any sensitive observations. Finally, our approach integrates predictions from pre-trained large language models. These models are known to be biased and capable of hallucinating and fabricating outputs. Spatially localizing the outputs of such models runs the risk of amplifying such biases if used inappropriately.
Appendix
In this appendix we provide additional quantitative and qualitative results, analysis, implementation details, and ablations.
In Appendix A, we provide additional results for uncertainty quantification, use of taxonomic rank text, a ‘distance weighted’ MAP metric, and show expanded results from Figure 3 for the non-low-shot setting. In Appendix B, we perform an ecologically relevant analysis of our results, showing how performance varies with region, range size, and taxonomic group. In Appendix C, we provide details on the implementation of FS-SINR and the baseline approaches, and of the training and evaluation procedure. In Appendix D, we provide additional ablations of FS-SINR, investigating the impact of training data, different input features, and modifications to the architecture. Finally, in Appendix E, we show additional qualitative results, including visualizing zero-shot and few-shot ranges for species and non-species concepts, and comparisons to ranges produced by LE-SINR and SINR approaches.
Appendix A Additional Quantitative Results
A.1 Uncertainty Quantification
Here, we report results for an ensemble based on FS-SINR and quantify the uncertainty in the ensemble’s predictions using methods adapted from Poggi et al. (2020). We create an ensemble by averaging the predictions of three FS-SINR models trained with different random seeds. Figure A25 shows range estimates from three such models, where we can see that each model can produce significantly different outputs. We take the average of these models as the ensemble prediction and treat the variance between individual model predictions as an estimate of the uncertainty of the ensemble. If all models agree that a species is either present or absent at a location, then the uncertainty will be low, while if models have different predictions, then the uncertainty will be high. in Table A1, comparing ‘Ensemble’ MAP to ‘Model’ MAP (repeated from FS-SINR ‘Text’ MAP in Table A4 for convenience), we can see that creating an ensemble increases MAP by 0.01 to 0.02, agreeing with typical findings that ensembling can improve performance relative to individual models (Dietterich, 2000).
In order to quantify uncertainty of our ensemble we follow an approach used in Poggi et al. (2020). We iteratively remove locations from the evaluation dataset that have the highest estimated uncertainty and recalculate the MAP without these locations. In our case we remove 2% of the data at each step. If the ensemble uncertainty aligns with how likely it is to be incorrect, then the MAP at each step will increase as the data with higher estimated uncertainty is removed from the evaluation. We can then plot the MAP against the fraction of data used for the evaluation. Taking the area under the curve generated doing this gives us the estimated ‘Sparsification Error AUC’ (SEAUC). If, instead, we remove 2% of locations randomly at each step then the MAP for each evaluation will remain approximately equal to the MAP when using the entire evaluation dataset. The ‘random’ SEAUC in this case is effectively equal numerically to the MAP using all evaluation data and can be estimated as such. Taking the difference between the estimated SEARC and the MAP then gives us the ‘Area Under the Random Gain’ (AURG). A positive AURG shows that the ensemble’s estimate of how uncertain its predictions are is better than random guessing. In Table A1 we see that the AURG is positive for all number of context locations and increases as more context locations are provided, showing that FS-SINR ensembles can provide a useful estimate of how certain they are about a prediction and that providing more context locations allows the ensemble to be more accurate in this uncertainty estimate.
In Figure A1 we visualize the ranges (means) and uncertainties (variances) for our FS-SINR ensemble for the Yellow-footed Green Pigeon, using ‘Range’ text, ‘Habitat’ text, or a single context location. We observe that the mean is high in the region of the expert-derived range, and lower in areas far from the range where a single model has erroneously predicted presence. The variance is generally lower in the region of the expert-derived range where all models agree, higher at the edges of this region where different models have different estimates of the extent of the range, and high in areas far from the true range where single models have incorrectly predicted presence, such as South America when ‘Habitat’ text is provided, or parts of Africa when ‘Range’ text is provided.
Range (Mean)
Uncertainty (Variance)
Range
Habitat
No Text
Figure A1: Range estimates and visualized uncertainty estimates for a FS-SINR ensemble. We display range estimates and uncertainties for the Yellow-footed Green Pigeon from an ensemble of three FS-SINR models. Zero-shot estimates are based on ‘Range’ text (top) and ‘Habitat’ text (middle). A few-shot estimate using no text and a single context location (bottom) is also shown. Range estimates for the ensemble (left) are a mean average of individual model predictions, while uncertainties (right) are estimated using the variances of the model predictions. The uncertainty is lower in the region of the expert-derived range where all models agree, higher at the edges of this region where different models have different estimates of the extent of the range, and high in areas far from the true range where single models have incorrectly predicted presence.
Table A1: Uncertainty quantification with ensembles. We show MAP on the S&T dataset for an ensemble of three FS-SINR models using ‘Range’ text and differing numbers of context locations (Ensemble MAP), with metrics for uncertainty quantification adapted from Poggi et al. (2020). We report ‘Sparsification Error AUC’ (SEAUC) and ‘Area Under the Random Gain’ (AURG) for the ensemble. Positive AURG shows the ensemble is performing better than random chance at estimating its uncertainty. We also present results for the same text and context locations for individual FS-SINR models for comparison (Model MAP). Ensembling slightly improves performance for all numbers of context locations.
# Context Model MAP Ensemble MAP SEAUC AURG
0 0.64 0.66 0.68 0.03
1 0.66 0.68 0.71 0.03
2 0.67 0.69 0.73 0.03
3 0.68 0.70 0.74 0.04
4 0.69 0.71 0.75 0.04
5 0.70 0.71 0.75 0.04
8 0.71 0.72 0.76 0.04
10 0.72 0.73 0.77 0.04
15 0.72 0.73 0.78 0.05
20 0.72 0.74 0.79 0.05
50 0.73 0.74 0.78 0.05
A.2 Taxonomic Understanding
[…]
Class
Aves
Family
Columbidae
Genus
Treron
Species
Phoenicopterus
[TRT Text]
[TRT Text] + 1 Context Location
Figure A2: Zero-shot and one-shot range estimation using Taxonomic Rank Text (TRT). Range predictions for the species Yellow-footed Green Pigeon from FS-SINR model trained on taxonomic rank text as in LD-SDM (Sastry et al., 2023), with expert-derived range inset. As seen in Figure A3 and Table A2, the text-based zero-shot predictions seem to more closely match the expert-derived range as more of the taxonomic rank text of the species is provided. Taxonomic rank text allows the model to somewhat localize predictions to areas where species sharing the provided taxonomy ranks are present in the training set. For example, Birds are globally distributed and we see the model attempt to output this in the zero-shot ‘Class’ visualization. Pigeons and Doves are not found in the extreme north and providing these ranks reduces predictions in these areas (and much of the northern hemisphere). The model mostly manages to identify that Green Pigeons are found only in Africa and parts of Asia. A single observation significantly contracts the predicted ranges, particularly when less taxonomic information is provided. Click on the taxonomic rank names to visit the iNaturalist page for that taxa where the geographic distribution of observations for it can be observed.
Here, we investigate the impact of providing FS-SINR with an understanding of the species taxonomy. For this we provide ‘Taxonomic Rank Text’ (TRT) instead of the Wikipedia-based free-form descriptions of a species that are used for our standard FS-SINR approach. This text gives the taxonomy of the species in decreasing taxonomic rank, in the form ‘class order family genus species’, so for a dog we would give the text ‘Mammalia Carnivora Canidae Canis Familiaris’. During training, we select a rank uniformly at random and remove all ranks beneath that. We hope that this process will force the model to learn an understanding of the distributions of not only individual species, but also genera, families, etc.. This may be helpful when facing unseen species as knowledge of the genus or family may provide clues about where this species may be found. This is similar to the approach used by LD-SDM (Sastry et al., 2023).
In Table A2 we show zero-shot performance for FS-SINR models trained on TRT on the IUCN and S&T evaluation tasks. We see that as we provide additional taxonomic information zero-shot performance improves, though it is still much worse than using habitat or range text. This implies that the model has managed to develop some understanding of the distributions of genera etc. and can use this to help map a novel species that shares higher order taxonomy with species in the training set.
In Figure A2 we provide some qualitative zero-shot and few-shot results showing the impact of training on taxonomic text. We see that the model appears to narrow down on the correct range as more specific taxonomy is revealed to it, from predicting across the entire globe when just the class Aves is provided, to removing northern latitudes as the family Columbidae is added, and finally removing the new world when the genus is provided. This broadly matches the actual distribution of these taxonomic ranks. Note that the relationship between taxonomic hierarchy and species range is likely complex as many speciation events (i.e., when a species splits into two or more new ones) can be the result of physical geographic barriers separating populations over time.
In Figure A3 we show few-shot results for FS-SINR models trained on TRT on the IUCN and SNT evaluation datasets. Zero-shot performance improvement with increasing taxonomic information is evident, but after very few provided locations this effect seems to disappear.
Method Variant IUCN S&T
FS-SINR 0.05 0.18
FS-SINR HT 0.33 0.53
FS-SINR RT 0.52 0.64
FS-SINR Class 0.05 0.19
FS-SINR Order 0.06 0.20
FS-SINR Family 0.12 0.25
FS-SINR Genus 0.18 0.30
FS-SINR Species 0.21 0.34
Figure A3: Impact of training and evaluating with Taxonomic Rank Text. Here we evaluate FS-SINR models trained using different context information on the IUCN dataset (left), and the S&T dataset (right). ‘Class’ indicates that only the taxonomic class of the species is provided as text during evaluation. ‘Order’ indicates that the taxonomic class followed by the order is provided as a text string during evaluation, and so on, such that ‘Species’ indicates that a text string in the format ‘class order family genus species’ is provided during evaluation. Providing more specific taxonomic text increases zero-shot performance. This is also presented Table A2. However we see that even the full taxonomy does not provide as much signal as habitat and range text for zero-shot range mapping. These more detailed texts provide more useful information for zero-shot range mapping - either actually mentioning geographic locations in the case of range text, or allowing the model to narrow predictions down to areas with specific features such as mountains and forests in the case of habitat text. When a single context location is provided, the choice of taxonomy text no longer seems to impact performance at all. It is possible that training on these less informative tokens means the model learns to pay less ‘attention’ to these text tokens compared to the Wikipedia-based text tokens usually used during training. This could explain why different rank taxonomy text tokens seemingly provide no benefit when any context locations are provided to the model.
A.3 Alternative Evaluation Metric
$w_{\bm{x}}=1+\frac{d_{range}(\bm{x})}{d_{antipodal}}{h},$ (3)
Results on the IUCN evaluation dataset can be found in Figure A4. We do not present results for the S&T dataset as we require knowledge of the range of each species globally to fairly apply the distance-weighted MAP, while the S&T dataset only provides range estimates for portions of the globe for each species. As the weight is increased, we observe a general reduction in overall performance. While there is no change in the relative ordering of different models, and FS-SINR outperforms LE-SINR and SINR across all settings of $h$, we do observe that FS-SINR and LE-SINR models that use habitat text during evaluation seem to decrease in performance more with larger $h$ compared to other approaches. They are likely most effected by the larger weight, as habitat text can cause the model to predict presence in locations around the world with similar habitat features such as mountains, forest, or desert, despite these locations being far from the true range. This appears to be true of both FS-SINR and LE-SINR. For LE-SINR, using habitat text outperforms not using text for the unweighted MAP, but using habitat text performs worse than not using text for the weighted MAP. In Figure A5, we display zero-shot results for two species where there is a large difference in performance based on the two metrics. In both cases FS-SINR using only text incorrectly predicts the species to be present far from the expert-derived range.
Unweighted MAP, $h=0$
Weighted MAP, $h=9$
Weighted MAP, $h=99$
Gravenhorst’s Mabuya African Jacana
[…]
Habitat Text: “T. gravenhorstii prefers rocky areas, grassland, shrubland, and forest, but is also found in coffee plantations and ylang-ylang plantations. The species is adapted to a variety of environmental conditions, including different types of ecosystems.” Habitat Text: “The African jacana prefers shallow lakes and its preferred habitat is floating vegetation. It has a specific environmental condition, which is not mentioned in the article, but it can be inferred that it requires a certain level of water depth and vegetation cover.”
A.4 Expanded Results
In Tables A3 and A4 we present and expanded set of results from Figure 3, for the IUCN and S&T datasets. We also report additional results for the non-few-shot setting using 500 and 1,000 context locations. Performance for most approaches including FS-SINR seems to plateau around 50 context locations, with some approaches gaining a minor boost in performance in the non-few-shot setting (SINR and Prototype-SINR), while others perform much worse (Active SINR). We also report results for a model trained and evaluated using context locations and features from a visual encoder with a DINOv2 backbone. This performs worse and has significantly more variable performance across runs compared to the pre-trained EVA-02 ViT, which was fine-tuned using species images during its original pre-training, used in the main paper.
Table A3: IUCN zero-shot and few-shot results. Here we present expanded IUCN evaluation results for the models shown in Figure 3 in tabular form. We also show results using a DINOv2 based image encoder. SINR and LE-SINR without text cannot produce a range map without at least one context location. Results are presented as MAP, where higher is better.
FS-SINR LE-SINR SINR Prototype SINR Active SINR
[…]
0 0.52 0.19 0.46 0.05 0.13 0.48 - - - -
1 0.57 0.45 0.55 0.48 0.40 0.55 0.47 0.42 0.48 0.48
2 0.60 0.54 0.59 0.56 0.49 0.57 0.52 0.47 0.53 0.55
3 0.62 0.58 0.61 0.60 0.53 0.58 0.54 0.50 0.56 0.58
4 0.63 0.60 0.62 0.62 0.55 0.59 0.56 0.52 0.57 0.59
5 0.64 0.62 0.63 0.63 0.56 0.60 0.57 0.54 0.58 0.60
8 0.65 0.64 0.64 0.65 0.59 0.61 0.59 0.56 0.60 0.62
10 0.66 0.65 0.65 0.66 0.60 0.62 0.60 0.57 0.61 0.62
15 0.67 0.66 0.66 0.67 0.61 0.63 0.62 0.59 0.61 0.62
20 0.67 0.66 0.66 0.67 0.61 0.64 0.63 0.61 0.62 0.62
50 0.68 0.67 0.67 0.67 0.61 0.66 0.66 0.64 0.62 0.60
500 0.68 0.66 0.66 0.67 0.57 0.67 0.67 0.65 0.63 0.37
1000 0.68 0.66 0.66 0.67 0.57 0.67 0.67 0.65 0.63 0.36
Table A4: S&T zero-shot and few-shot results. Here we present expanded S&T evaluation results for the models shown in Figure 3 in tabular form. We also show results using a DINOv2 based image encoder. SINR and LE-SINR without text cannot produce a range map without at least one context location. Results are presented as MAP, where higher is better.
FS-SINR LE-SINR SINR Prototype SINR Active SINR
[…]
0 0.64 0.38 0.64 0.18 0.28 0.60 - - - -
1 0.66 0.49 0.66 0.50 0.44 0.64 0.52 0.49 0.54 0.53
2 0.67 0.57 0.67 0.58 0.53 0.66 0.57 0.55 0.59 0.59
3 0.68 0.61 0.68 0.61 0.57 0.67 0.60 0.58 0.61 0.62
4 0.69 0.64 0.69 0.64 0.61 0.67 0.61 0.59 0.62 0.65
5 0.70 0.66 0.70 0.65 0.63 0.68 0.62 0.60 0.63 0.65
8 0.71 0.69 0.71 0.68 0.65 0.69 0.65 0.63 0.64 0.67
10 0.72 0.70 0.71 0.69 0.67 0.69 0.66 0.64 0.65 0.67
15 0.72 0.71 0.72 0.70 0.68 0.70 0.68 0.67 0.65 0.67
20 0.72 0.71 0.72 0.71 0.68 0.71 0.69 0.68 0.65 0.66
50 0.73 0.72 0.73 0.71 0.69 0.73 0.72 0.72 0.66 0.64
500 0.73 0.71 0.72 0.71 0.62 0.73 0.72 0.73 0.67 0.24
1000 0.72 0.71 0.72 0.71 0.62 0.73 0.72 0.73 0.67 0.24
Figure A6: Comparison to traditional machine-learning approaches. Here we evaluate a Gaussian Process approach and a Random Forest on the IUCN (left) and S&T (right) datasets. We find that both approaches perform far worse than FS-SINR, LE-SINR or SINR for species range estimation using presence-only data. ‘Range’ text is used here for FS-SINR and LE-SINR, while other approaches take only context locations as input.
A.5 Comparison to Traditional Machine Learning Methods
In Figure A6 we compare FS-SINR with two traditional machine learning approaches for classifying whether a species is present or absent at a given location. Both approaches were implemented using scikit-learn (Pedregosa et al., 2011). We compare with a Gaussian Process approach adapted from Golding & Purse (2016). While this method is designed for presence-absence data, we adapt it for the presence-only setting by providing a number of pseudo-negatives equal to the provided number of context locations, which performed best out of several strategies we investigated. Using a large number of pseudo-negatives as we do for FS-SINR both performed poorly and took excessively long to run as Gaussian Process computations scale cubically with the amount of data. For this approach we use a logit link function and a squared-exponential kernel. We also compare to a random forest classifier (Breiman, 2001). We investigated providing the same number of pseudo-negatives as for FS-SINR with appropriate class weights, however we found better performance by providing the same number of pseudo-negatives as context locations. In both cases, performance for this task is significantly worse than FS-SINR, LE-SINR and SINR.
Appendix B Ecologically Relevant Analysis of Results
To give a more ecologically relevant analysis of our results we present them here by continent, species range size, and taxonomic class. In these cases we display results for the IUCN evaluation dataset only, as it provides expert-derived presence-absence information globally and includes a range of taxonomic groups. In comparison the S&T dataset only includes Aves, i.e., bird species, and evaluates the presence or absence of each species over a portion of the globe, preventing us from calculating global range sizes and performance by continent.
B.1 Results by Region
In Figure A7 we show performance of FS-SINR, LE-SINR, and SINR models by continent for few-shot and zero-shot text-only predictions. FS-SINR outperforms other approaches on all continents except South America and Oceania, where at larger numbers of context locations LE-SINR becomes comparable. Biodiversity data and particularly global-scale citizen science datasets such as the iNaturalist-derived data we use to train FS-SINR can contain large biases (Geldmann et al., 2016; Hughes et al., 2021), and here we can see the impact of this bias within our training data.
We have more species observation training data (taken from Cole et al. (2023), which also visualizes the data distribution) from Europe and North America than from other areas, and the observations in these regions are relatively well distributed spatially. In other continents, there are large areas with very few or no training observations for models to learn from, along with small areas that are highly observed. On top of this, our text descriptions are taken from English language Wikipedia and may be more descriptive for species found in areas where English is widely spoken, and our pre-trained large language model may have more knowledge of North American and European geography and ecology due to biases in the text data used for training. Combined, these factors lead to higher performance in North America and Europe compared to other regions. In Figure A10 we show the average false positive error for few-shot range estimation for FS-SINR on the IUCN evaluation dataset. This indicates greater error in regions where we have less training data.
B.2 Results by Species Range Size
Here, we display results showing the average MAP for species in our IUCN evaluation dataset, grouped by range size, where range size is computed from the expert-derived range maps. In Figure A9 we break down performance of zero-shot approaches by range size for FS-SINR, LE-SINR, and SINR. We find that for all models and settings, performance varies very strongly with range size. This is most significant in the zero-shot setting. FS-SINR performs well compared to the baselines, though all models struggle with very small ranges. As small-ranged species are especially vulnerable to extinction (Chichorro et al., 2019), current methods performing poorly for these species when evaluated globally may be of concern and improving the modeling of these species may be a priority from a conservation perspective. We also see that performance worsens for the very largest ranges.
B.3 Results by Taxonomic Class
Here we break down performance on the IUCN evaluation dataset by taxonomic class. Four taxonomic classes are present in our training data, namely Amphibia, Aves, Mammalia, and Reptilia. In Figure A8 (a) we display zero-shot performance for FS-SINR and LE-SINR using range and habitat text. We observe that Aves and especially Mammalia outperform the other classes, particularly when habitat text is provided. Albert et al. (2018) suggest that of the 20 most ‘charismatic’ species in the western world, all but the Great White Shark and Crocodile are mammals, and Trimble & Van Aarde (2010) show that scientific research is heavily focused on mammals. We may be seeing the impact of this, where mammals are more likely to have detailed Wikipedia pages which we draw our textual training and evaluation data from. In Figure A8 (b), (c), and (d), we investigate how these differences in performance between taxonomic classes change as more location data is provided. We see that for both FS-SINR and LE-SINR, providing context locations significantly reduces the differences in performance between taxonomic class, though mammals do continue to very slightly outperform other taxonomic groups for a given model and setting.
(a) 0 Context Locations
(b) 5 Context Locations
(c) 20 Context Locations
(d) 50 Context Locations
(a) 0 Context Locations
(b) 5 Context Locations
(c) 20 Context Locations
(d) 50 Context Locations
(a) 0 Context Locations
(b) 5 Context Locations
(c) 20 Context Locations
(d) 50 Context Locations
Range Text
Habitat Text
No text
1 Context Location
5 Context Locations
20 Context Locations
50 Context Locations
Figure A10: Average false positive error by location for few-shot approaches. Here we see average false positive error of FS-SINR on the IUCN dataset. Providing any text leads to an increase in the false positive error, although Figure 3 suggests that this text still helps with range estimation performance. As the number of provided context locations increases, the impact of the text is reduced and the distribution of errors appear similar.
Appendix C Implementation Details
C.1 FS-SINR
C.1.1 Model Architecture
@@
Code for FS-SINR is available at https://github.com/Chris-lange/fs-sinr. The architecture for FS-SINR consists of four components: the location encoder, $f$; the text encoder, $t$; the image encoder, $a$; the transformer encoder, $m$; and the species decoder, $s$. These components comprise 8,154,368 learnable parameters in total. All non-linearities in FS-SINR are ReLUs.
@@
The location encoder, $f$, is identical to the one used in Hamilton et al. (2024), which is taken from  Cole et al. (2023). It is composed of an initial linear layer and ReLU non-linearity followed by four residual layers, where each is a two-layer fully connected network with residual connections (He et al., 2015) between the input and output of each residual layer. Each layer contains 256 neurons, and there are 527,616 learnable parameters in total.
@@
The text encoder, $t$, follows the structure of the text-based species encoder from Hamilton et al. (2024). In $t$, a pre-trained and frozen large language model, GritLM (Muennighoff et al., 2025), is used to produce a fixed 4,096 length embedding from input text. This is then passed through a smaller network to reduce the dimensionality to 256. This smaller network consists of two residual layers with a hidden layer size of 512. In total, the text encoder contains 3,410,432 learnable parameters.
@@
The image encoder, $a$, has a structure similar to $t$. In $a$, a pre-trained and frozen vision transformer, EVA-02 (Fang et al., 2024), pre-trained on images from 10,000 species from the iNaturalist species classification dataset (Van Horn et al., 2021), is used to produce a fixed 1,024 length embedding from an input image, by extracting the CLS token from the final layer of the model. This is then passed through a smaller network to reduce the dimensionality to 256. This smaller network consists of two residual layers with a hidden layer size of 512. In total, the image encoder contains 1,837,568 learnable parameters. Tables A3, A4 and A5 include results using a DINOv2-large image encoder (Oquab et al., 2024) instead of the EVA-02 ViT where all other architecture choices remain the same.
@@
The transformer encoder, $m$, takes in an arbitrary length set of unordered 256 dimensional tokens produced by $f$, $t$, and $a$, as well as two learned tokens that are added to each set of inputs. The CLS, class, token produces the species range, and a ‘Register’ token, inspired by Darcet et al. (2024), which acts as an additional repository of global information during encoding. Element-wise addition between each token and one of five learned 256 dimensional ‘token type embeddings’ is performed to allow the model to differentiate between tokens from different sources. The transformer itself is composed of four transformer encoder layers, implemented using PyTorch’s nn.TransformerEncoderLayer (Paszke et al., 2019), based on Vaswani et al. (2017). Key-query-value multi-head attention is used with two ‘heads.’ The feed forward components contain 512 neurons per layer, while the token dimensionality is 256. Layer norm is used in each layer, using a default epsilon value of 1e-5 for enhanced numerical stability. In total, $m$ contains 2,176,256 learnable parameters. Finally the species decoder, $s$, is a simple fully connected network with two hidden layers. Each layer contains 256 neurons, and in total the decoder contains 197,376 learnable parameters.
C.1.2 Training